OpenStudy (anonymous):

The width of a rectangle is half as long as the length. The rectangle has an area of 200 square feet. What are the Length and width of the rectangle?

OpenStudy (id21):

width is x=10, length is 2x= 20

OpenStudy (id21):

Area is \[2x ^{2}=200\]

OpenStudy (anonymous):

Let the sides be a and b. a = 1/2b; area = a*b = 200 sq.ft. substitute a and get 1/2 b*b = 200 1/2 b^2 = 200 b^2 = 400 b = 20 so a = 1/2 b = 10.